OpenStudy (anonymous):

What is the remainder of (x^3 – 6x – 9x + 3) ÷ (x - 3)

OpenStudy (anonymous):

is it -6x or 6x^2?

OpenStudy (anonymous):

-6x

OpenStudy (anonymous):

and 9x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you can find that out by either doing euclidean division of polynomials. A general way to do euclidean division of polynomials is polynomial long division. A specific (and faster) way to do it for cases where the denominator is in the form of (x-a), is to use synthetic division.

OpenStudy (anonymous):

then eq is x^3 -15x +3 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

remainder is -15

OpenStudy (anonymous):

Thats not an answer :/ A. -51 B. -51/x-3 C. -17 /x D. -17 /x-3

OpenStudy (anonymous):

the answere is -15

OpenStudy (anonymous):

check your problem once more, is it right?? see your book

OpenStudy (anonymous):

@jimmyao98

OpenStudy (anonymous):

I'll take a screenshot

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

e) none of above...

OpenStudy (lgbasallote):

you can just use remainder theorem...

OpenStudy (anonymous):

-15 maybe noanswer

OpenStudy (lgbasallote):

P(3) = (3)^3 - 6(3) - 9(3) + 3 = 27 - 18 - 27 + 3 = -18 + 3 = -15

OpenStudy (anonymous):

f) teacher fail

OpenStudy (lgbasallote):

computer fail*

OpenStudy (anonymous):

g) not understanding failure of human input fail.

OpenStudy (anonymous):

clearly supposed to be -6x^2 go point out the mistake to your teacher and tell him you deserve bonus credit.

OpenStudy (anonymous):

tell him @Algebraic! said it was ok. He'll know what you mean.

OpenStudy (anonymous):

h) kill teacher

OpenStudy (anonymous):

I actually thought that it was that mistake and put in -51 and still got it wrong